### v1

# How spatial patterns can lead to less resilient ecosystems

Abstract
Several theoretical models predict that spatial patterning increases ecosystem resilience. However, these predictions rely on strong simplifying assumptions, such as isotropic and infinite ecosystems, and we lack empirical evidence directly linking spatial patterning to enhanced resilience. We introduce a unifying framework, encompassing existing models for vegetation pattern formation in water-stressed ecosystems, that relaxes these assumptions. This framework incorporates finite vegetated areas surrounded by desert and anisotropic environmental conditions that induce non-reciprocal plant interactions. Under these more realistic conditions, we identify a novel desertification mechanism, known as convective instability in physics but largely overlooked in ecology. These instabilities form when non-reciprocal interactions destabilize the vegetation–desert interface and can trigger desertification fronts even under stress levels where isotropic models predict stability. Importantly, ecosystems with periodic vegetation patterns are more vulnerable to convective instabilities than those with homogeneous vegetation, suggesting that spatial patterning may reduce, rather than enhance, resilience. These findings challenge the view of self-organized patterns as indicators of resilience and provide a new framework to investigate how spatial dynamics determine the stability and resilience of ecological systems across scales.
1 Introduction
Ecosystems can exhibit multiple alternative stable states, each with different structures and functions [1, 2]. Due to the nonlinear processes underlying ecosystem dynamics, small changes in environmental conditions can trigger persistent changes between these states once a critical threshold, or tipping point, is crossed [3]. These transitions, known as regime shifts, may propagate over large distances and even different biomes [4], impacting ecosystem services and human well-being substantially [5, 6]. Given the severe consequences of regime shifts and the challenges in predicting or reversing them, much research has focused on developing predictive theories of regime shifts as well as quantitative indicators to anticipate and potentially mitigate their effects [7, 8, 9].
Regime shifts have been observed in different terrestrial and marine ecosystems, such as shallow lakes, savannas, kelp forests, or drylands [10]. Dryland regime shifts occur when aridity crosses a critical threshold, causing vegetation loss and the ecosystem collapse into a desert state [11]. Models suggest that spatial processes—particularly self-organized regular patterns of vegetation and bare soil—enhance dryland resilience and resistance to external perturbations [12, 13]. First, these patterns emerge once a resource scarcity threshold is crossed and adapt their shape as resource availability diminishes further [14, 15], suggesting patterns could be a signature of stress and indicate an ongoing desertification process and proximity to a full vegetation collapse [16]. Second, ecosystems exhibiting periodic patterns can survive beyond the tipping point predicted by non-spatial theories, thereby enhancing their resistance at higher aridity levels [17]. Finally, patterns with slightly different spatial properties can coexist within a range of environmental conditions, suggesting that patterns are an adaptive feature allowing ecosystems to buffer external perturbations [13, 18, 19]. However, due to the large scales involved in pattern formation and dryland vegetation dynamics, the possibilities for testing these predictions are limited [18, 20]. Consequently, whether and how spatial patterning impacts ecosystem resilience remains unknown [21, 22].
The relationship between vegetation patterns and desertification processes has been mainly studied using simple models for flat, isotropic landscapes and infinitely large vegetated areas. However, real vegetation patterns cover large, but finite, regions [23], and are often embedded in landscapes with different topographies, such as hillsides or microreliefs [24, 25, 26, 27]. Topography breaks spatial isotropy, making interactions between vegetation patches depend not only on their distance, but also on their relative position, i.e., non-reciprocal. More importantly, assuming infinite landscapes oversimplifies several spatial processes that can cause vegetation loss and tipping, the most important being the propagation or reversal of desertification fronts in response to increased aridity [28, 29]. Desertification front propagation can occur prior to tipping points, turning abrupt regime shifts into gradual [30], and is sensitive to boundaries, spatial heterogeneities, and the size of the vegetated area. Spatial heterogeneity in plant interactions and boundary effects have been separately included in studies of vegetation pattern formation to explain the formation of vegetation stripes [31, 32] or the spreading of invasive species [33], among other phenomena. Yet, only recently have studies included both to successfully explain topological properties observed in banded vegetation patterns worldwide [34]. How much they influence the stability of vegetated states in drylands remains to be explored.
In this study, we address this gap by investigating the effects of spatial heterogeneity and non-reciprocity, induced by boundary effects and environmental anisotropies, on desertification dynamics. Our results show that non-reciprocity in plant interactions enhances the invasion of desertification fronts into vegetated areas, thereby increasing the likelihood of regime shifts at lower environmental stress. Moreover, this phenomenon is more pronounced in the presence of vegetation patterns, causing self-organized ecosystems to collapse at a stress level at which non-patterned vegetation survives. These findings suggest that spatial self-organization may increase the susceptibility of ecosystems to regime shifts and, consequently, reduce their resistance to environmental change, thereby questioning the current consensus regarding the role of self-organized patterns as indicators of ecosystem resilience.
2 Mathematical modeling
Several models have been proposed to describe vegetation dynamics in water-limited ecosystems. We will focus on models describing vegetation as a continuous biomass density $b(x,t)$ evolving in time and space according to partial differential equations (PDE), which is the usual choice to study regular vegetation patterns [35]. The diversity of PDE-based models for vegetation dynamics can be organized into two main categories: nonlocal interaction redistribution models, where plant interactions are described implicitly by kernel functions that modulate the intensity of positive and negative density-dependence in plant growth, and Turing-like reaction-diffusion models describing water-vegetation feedbacks explicitly [36]. Although they are mathematically different, models within these two categories can be reduced to a general equation near the onset of instability of the unpopulated state (Fig. 1). We exploit this universality and perform our analyses using this reduced equation, which makes our results independent of specific modeling assumptions. We provide a full derivation of this equation in the Supplementary Material A, starting from two prototypical examples of kernel-based and reaction-diffusion models, while we focus here on presenting the reduced model and discussing its general structure. In its dimensionless form, this equation reads
$\frac{\partial b}{\partial t}$ $=$ $-\eta b+\kappa b^{2}-b^{3}/2+d\frac{\partial^{2}b}{\partial x^{2}}$ (2.1)
$-b\left(\alpha\frac{\partial}{\partial x}+\Gamma\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{4}}{\partial x^{4}}\right)b,$
where the specific ecological meaning of each parameter will depend on which model we use to derive it (see Supplementary Material and Fig. 1). In general, Eq. (2.1) consists of a linear net death term $-\eta b$, where $\eta$ can be negative to represent growth, and local competition and facilitation contained in the terms $\kappa b^{2}-b^{3}/2$. Ecologically, these first three terms define a nonlinear density-dependent growth rate due to, for example, an Allee effect. The spatial dynamics is described by a linear diffusion accounting for plant dispersal with diffusion constant $d$, and a nonlinear spatial differential operator. In this nonlinear operator, the term proportional to $\alpha$ originates from isotropy-breaking processes while the last two terms encapsulate the pattern-forming feedbacks. Specifically, the combination of the linear and density-dependent nonlinear diffusion leads to an effective diffusion term, $d-b\Gamma$, that can become negative and trigger spontaneous aggregation of the biomass for sufficiently large $\Gamma$. This aggregation process is saturated by the last term, resulting in stable patterns. Three parameters are particularly relevant for our analysis. First, the net growth rate $-\eta$ measures the difference between baseline death and growth rates and hence serves as a proxy for environmental stress. Large positive values of $\eta$ represent high stress, and stress decreases as $\eta$ decreases, eventually becoming negative when environmental conditions are such that they can sustain baseline population growth. Second, $\alpha$ modulates the only term breaking the isotropy in Eq. (2.1) and hence controls the intensity of nonreciprocity in plant-plant interactions. This symmetry-breaking can be caused by various factors, such as slopes, more complex but uni-directional topographies, or fog water carried by wind [32, 27, 37]. Finally, the parameter $\Gamma$ directly controls the pattern formation, allowing us to study homogeneous or patterned ecosystems by varying this single parameter.
Figure 1: Schematic summary of model reduction from two representative examples of reaction-diffusion and nonlocal interaction-redistribution models to the reduced equation (2.1). The parameters of the reduced equation encapsulate the ecological feedbacks in each of the original models as described in the table.
Finally, to fully describe the model, we need to specify the boundary conditions for Eq. (2.1). Most studies often assume periodic boundary conditions, mimicking infinitely large and spatially invariant ecosystems. This assumption, while mathematically convenient, is not realistic in most ecosystems, where vegetation (patterned or not) is often embedded in non-vegetated areas or at the edge of a dryland desert transition [14, 23]. To mimic these conditions, we consider boundary conditions such that vegetation biomass vanishes at the system edges $x\leq 0$ and $x\geq L$ (with $L$ the system length). These boundary conditions imply that any solution in the ecosystem bulk has to connect with the unvegetated, $b=0$ solution at the boundaries, which is also a solution of Eq. (2.1). Thus, a front naturally exists for all the solutions except the unvegetated one, which is the only strictly homogeneous solution possible in the system. The motion of this front will dictate whether the desert state $b=0$ propagates into the vegetated region or not. This motion is directly determined by the parameters of Eq. (2.1), which account for the strength of the different plant interactions and environmental stresses. We next analyze this front propagation process, both in cases where plant biomass in the vegetated area is quasi-homogeneous (i.e., homogeneous within the bulk) or forms self-organized regular patterns.
3 Results
3.1 Stationary solutions
Due to boundary effects, where the vegetation density must go smoothly to zero, our model does not allow for strictly homogeneous states with $b\neq 0$. However, far from these vegetation-desert boundaries (i.e., in the bulk of the vegetated area), we can disregard boundary effects and obtain quasi-homogeneous solutions from the nonspatial terms of Eq. (2.1). These solutions are $b_{\pm}(\eta,\kappa)=\kappa\pm\sqrt{\kappa^{2}-2\eta}$ and $b_{0}=0$, where $\kappa$ is related to facilitative interactions and controls the existence of alternative stable states and the linear mortality rate $\eta$ accounts for the intensity of environmental stress. For $\kappa<0$ (i.e., no net facilitation), the nonspatial model is monostable. If additionally $\eta<0$ (low stress), then $b_{0}$ is unstable, and $b_{+}$ is stable. At higher stress $\eta>0$, $b_{0}$ is the only possible steady state and thus stable. In the presence of facilitation, $\kappa>0$, the system presents two alternative stable states for stress levels $0<\eta<\eta_{c}$, with $\eta_{c}=\kappa^{2}/2$. $\eta_{c}$ is hence a tipping point such that environmental conditions worsening beyond this point induce a sudden regime shift from a vegetated $b_{+}$ to a desert $b_{0}$. Homogeneous solutions in the bulk $b_{+}$, connect with the boundary condition $b=0$ near the system edges. This solution naturally generates two interfaces, or fronts, in the system, one in the proximity of $x=0$ and the other in the proximity of $x=L$. As the parameter $\eta$ is controlled by the effective mortality of the plants (considering the environmental conditions), it is generally considered the control parameter.
Non-homogeneous solutions are much harder to obtain analytically, but we can compute the region of the parameter space where they occur by performing a linear stability analysis on Eq. (2.1). Again, our analysis focuses on the ecosystem bulk and therefore considers that the boundaries are infinitely far. To perform the linear stability analysis, we introduce a small perturbation to the uniform solution, $b=b_{+}+u_{0}e^{\lambda t+ikx}$ and linearize in $u_{0}$, which gives a perturbation growth rate of the form
$\lambda(k)=-\eta+2\kappa b_{+}-3b_{+}^{2}/2-(d-b_{+}\Gamma)k^{2}-b_{+}k^{4}-ib_{+}\alpha k.$ (3.1)
A heterogeneous solution may appear whenever $\lambda(k)>0$ for some $k\neq 0$, meaning that the term $\propto e^{ikx}$ will start to grow exponentially. This condition is true when the environmental stress is higher than a threshold value $\eta_{T}$, which we can obtain by simultaneously solving
$k_{c}^{2}$ $=$ $(b_{T}\Gamma-d)/(2b_{T}),$
00 0 $=$ $-\eta_{T}+2\kappa b_{T}-3b_{T}^{2}/2+b_{T}k_{c}^{4},$ (3.2)
with $b_{T}=b_{+}(\eta_{T},\kappa)$. This analysis allows us to identify $\Gamma$ and $d$ as the control parameters triggering a Turing instability and thus patterns. In the following sections, we investigate numerically the stability of patterned and quasi-homogeneous vegetation distributions in response to worsening environmental conditions (see Supplementary Material B for a description of the numerical method). We perform these analyses using the reduced Eq. (2.1), but our results also hold working with the original kernel-based and Turing-like models (see Supplementary Material C).
3.2 Desertification dynamics
To analyze the persistence of vegetated states under worsening environmental conditions and ecosystem tipping, we first fixed $\kappa>0$ to ensure bistability between a homogeneous cover and bare soil, $b_{+}$ and $b_{0}$ respectively. Using the results from the linear stability analysis performed in the previous section, we kept all parameters constant and varied $\Gamma$ to move between quasi-homogeneous and patterned solutions–low and high $\Gamma$, respectively–and $\alpha$ to compare ecosystems driven by reciprocal and non-reciprocal plant interactions.
3.2.1 Desertification by vegetation destabilization: ecosystem tipping.
In the simplest case with $\alpha=0$, because we choose a large enough vegetation patch, both patterned and non-patterned vegetation covers collapse when the environmental stress is large enough to induce vegetation loss everywhere in the patch. For quasi-homogeneous vegetation covers, the total biomass largely follows the analytical results obtained with the non-spatial model (brown dots in Fig. 2a), and the ecosystem undergoes a sudden desertification process once the environmental stress crosses the threshold $\eta_{c}$. The emergence of self-organized vegetation patterns makes the ecosystem more resistant to worsening environmental conditions, allowing vegetation to survive up to an environmental stress threshold $\eta_{pc}>\eta_{c}$ (brown dots in Fig. 2b). This result corresponds to the classical understanding of spatial self-organization in drylands [15, 38, 17, 13]. Notice, however, that we obtained this result by varying quasi-statically the environmental stress control parameter, such that the solution branches in Fig. 2a, b (brown dots) represent the observed equilibria of the system. More complex ways of varying the control parameter, such as parameter quenches or temporal variations, as well as intricate initial conditions, would lead to a much more complex diagram of states [39, 40, 30, 41, 42, 43, 23]. In addition to this desertification transition via global destabilization of the vegetation cover, environmental stress can also destabilize the two fronts at the vegetation-desert interfaces $x=0$ and $x=L$ and trigger a desertification front propagation. In principle, this interface destabilization occurs at environmental stress values below the tipping point, when $\eta$ crosses what is called a Maxwell point [44, 40]. However, we do not observe desertification by front propagation in this isotropic regime, probably because the Maxwell and the tipping point are very close to each other for all the parameter sets we considered.
Figure 2: a,b) Bifurcation diagrams showing the stability of the vegetation cover for varying levels of environmental stress, $\eta$ in a non-pattern (a) and pattern-forming (b; $\Gamma=2.1$) ecosystem. $b_{max}$ shows the maximum of the vegetation biomass in the ecosystem bulk (100 simulation points away from the boundary). Brown (dark) dots are obtained in isotropic conditions, $\alpha=0$, and orange (light) ones in relatively high anisotropic conditions. $\eta_{conv}$ indicates the new bifurcation points induced by nonreciprocity at which the vegetation cover collapses. c) In the $(\eta,\alpha)$ parameter space, the curves $\eta_{conv}(\alpha)$—dashed for patterned and solid for non-patterned ecosystems—delineate the boundaries between different regimes, each characterized by spatial configurations of vegetation cover that support ecosystem persistence. The critical value $\alpha_{c}$ marks the threshold of non-reciprocity beyond which patterned ecosystems become less resilient than their non-patterned counterparts.
3.2.2 Desertification by front propagation: convective instabilities.
Boundary effects become important, even for large systems, when considering non-reciprocal plant interactions, $\alpha\neq 0$. We consider only $\alpha>0$, which is equivalent to $\alpha<0$ under a $x\rightarrow-x$ reflection of the spatial coordinate. Regardless of whether the vegetation cover is quasi-homogeneous or forms spatial patterns, introducing non-reciprocal interactions anticipates the ecosystem collapse to a lower stress threshold $\eta_{conv}<\eta_{(c,pc)}$ (orange vs. brown dots in Fig. 2a,b). This decrease in ecosystem resistance to worsening environmental conditions results from a new desertification process driven by a convective instability at the vegetation-desert interface. For $\alpha=0$, this interface is stable because the desert boundary conditions prevent the expansion of the vegetation patch. However, the velocity of this front, including its direction, depends on model parameters, including $\alpha$. Therefore, increasing $\alpha$ eventually reverses the front velocity, triggering the propagation of a desertification front into the vegetation patch at a value of $\eta$ at which reciprocal ecosystems are stable. We use the shift in the stress threshold causing vegetation loss to quantify this loss in ecosystem resistance, $\Delta\eta_{(c,pc)}(\alpha)=\eta_{(c,pc)}-\eta_{conv}(\alpha)$, where the subscripts $(p,pc)$ indicate whether $\Delta\eta$ is calculated for quasi-homogenous or patterned biomass distributions, respectively.
To better understand how convective instabilities destabilize vegetated states, we thoroughly investigated how non-reciprocity in plant interactions and environmental stress jointly cause ecosystem collapse (Fig. 2c). In general, increased non-reciprocity in plant interactions leads to earlier ecosystem collapses regardless of the spatial distribution of vegetation (patterns vs. homogeneous). For weak reciprocity, patterned ecosystems are more resistant than non-patterned ones. However, because $\eta_{conv}$ decays with increasing $\alpha$ much faster for patterned ecosystems relative to non-patterned ones, there is a crossover between both curves at a critical strength of the non-reciprocity parameter $\alpha_{c}$. Remarkably, past this value $\alpha_{c}$, ecosystems exhibiting spatial patterns collapse at a lower environmental stress relative to those where plant biomass is homogeneously distributed within the ecosystem bulk. This result suggests that spatial self-organization could hinder ecosystem resistance to worsening environmental conditions, challenging the current consensus regarding the role of spatial self-organization as an enhancer of ecosystem robustness and resilience.
Finally, we measured how much and in which conditions spatial dynamics impact the environmental stress threshold triggering ecosystem collapse. We computed $\eta_{conv}$ for several different pattern wavelengths and amplitudes–mainly controlled by $\Gamma$–and intensities of non-reciprocity in plant interactions $\alpha$. Then, we compared it to the tipping point of the non-spatial system $\eta_{c}$. For weakly non-reciprocal interactions, $\eta_{conv}>\eta_{c}$, indicating that self-organized patterns enhance ecosystem resistance (top left region of Fig. 3). However, past a limit value of the non-reciprocity parameter (black dashed line in Fig. 3), $\eta_{conv}<\eta_{c}$, indicating that spatial effects reduce ecosystem resistance to worsening environmental conditions. Moreover, within this region of the parameter space, $\eta_{conv}$ decreases with increasing $\Gamma$ at constant $\alpha$. This behavior of the environmental stress threshold triggering convective instabilities indicates that ecosystems become weaker the more we enhance spatial pattern formation while keeping non-reciprocity intensity constant.
Figure 3: Behavior of the convective tipping point $\eta_{conv}$ as a function of the parameters $\Gamma$ and $\alpha$. The color map shows the value of $\eta_{conv}$ relative to $\eta_{c}$ (the tipping point of homogeneous systems), and a dashed line indicates the limit $\eta_{conv}=\eta_{c}$, separating regions in which patterns are more or less resilient than a corresponding homogeneous system (without spatial effects). Increasing $\Gamma$ triggers the Turing instability and changes the pattern morphology (wavelength and amplitude). Increasing $\alpha$ always destabilizes the ecosystem and its effect is more severe on patterned ones.
4 Discussion
Using a reduced equation representative of a broad variety of vegetation models, we analyzed spatiotemporal vegetation dynamics in confined environments where environmental anisotropies lead to non-reciprocal interactions between plants. These two features—spatial confinement and directional environmental forcing—are common in drylands, particularly across ecotones between (semi)-arid vegetated areas and deserts, where self-organized patterns are more common [45, 14]. Incorporating them into dryland models challenges the prevailing interpretation of self-organized vegetation patterns as indicators of enhanced ecosystem resilience [17, 21, 22, 46]. This established interpretation relies on theoretical results obtained under highly idealized conditions, assuming isotropic and infinitely large ecosystems [21, 36]. However, under the more realistic conditions we considered here, the vegetation cover becomes susceptible to convective instabilities, a well-known phenomenon in physical systems but, to the best of our knowledge, largely overlooked in ecological contexts. Convective instabilities destabilize the vegetation-desert interface, facilitating the propagation of a desertification front into vegetated regions and causing ecosystem collapses at lower levels of environmental stress than those predicted by idealized models. Importantly, patterns amplify this shift, causing self-organized ecosystems to collapse under stress levels that uniform vegetation can withstand. See Fig. 4 for a summary of our results.
Desertification by front propagation can induce gradual regime shifts in isotropic ecosystems, both for patterned and homogeneous vegetation [30, 47, 40]. However, spatial patterns result in more resilient ecosystems by slowing down—and in some cases even stopping [48]—front propagation. Moreover, the multistability of different wavelength patterns creates multiple thresholds (one for each wavelength) that need to be crossed to cause a complete ecosystem collapse, further enhancing the resilience of patterned isotropic ecosystems [49]. Our results indicate that the Maxwell point, at which the vegetation-desert interface destabilizes, is very close to the tipping point of non-spatial models, which does not allow us to observe desertification fronts for reciprocal plan interactions [28]. Yet, convective instabilities become evident even at apparently negligible levels of non-reciprocity, which suggests that this new desertification mechanism operates at significantly lower levels of environmental stress than both isotropic desertification fronts and abrupt tipping. Therefore, stable ecosystems under increasing environmental stress are likely to first exhibit sensitivity to convective instabilities. This highlights the importance of incorporating fine-scale environmental anisotropies—and their effects on plant-plant interactions—into dryland models. In simplified scenarios, such as those considered by our model, environmental anisotropies affect front dynamics but do not have any impact on total biomass and spatial pattern formation. In such cases, single-time snapshots of vegetation patterns may fail to reveal the presence of underlying environmental anisotropies unless the patterns themselves are clearly anisotropic [32, 34]. Consequently, evaluating desertification risk due to fronts driven by non-reciprocal interactions requires spatiotemporal datasets or direct field measurements [50]. In more complex setups, environmental anisotropies can be inferred by jointly analyzing remotely sensed vegetation data and other ecosystem features, such as topography [27, 51, 25], wind, and directional fog [52, 37]. Models accounting for such environmental complexity at these different scales will provide more reliable insights into how vegetation spatial dynamics influence ecosystem resilience [53, 54].
Figure 4: Schematic summary of the effect of non-reciprocal plant interactions in homogeneous (a) and patterned (b) ecosystems. Convective instabilities make ecosystems, with and without patterns, collapse before their respective isotropic tipping point $\eta_{conv}<\eta_{(pc,c)}$, as well as before their corresponding Maxwell point $\eta_{M}$. This shift is stronger for patterned vegetation, which can make self-organized ecosystems collapse before those with homogeneous vegetation. In terms of the vegetation-desert from velocity, we distinguish three regimes. At low environmental stress (green region in a, b; $\eta<0$), the $b=0$ is unstable and a vegetation front invades the desert region (leftmost panels in the bottom row of a and b). The desert state becomes stable at intermediate stress levels (yellow region in a, b), however, vegetation is still able to invade the desert state. Finally, the front velocity reverts when environmental stress increases (red region in a, b), triggering a desertification front propagation into the vegetated area, exacerbated by nonreciprocal interactions.
Another feature that makes desertification via convective instabilities particularly relevant is that they arise within broad parameter regimes, including weak non-reciprocal plant interactions, and for various modeling frameworks. Additionally, patterned ecosystems only require weak non-reciprocity to become more sensitive to convective instabilities than homogeneous vegetation covers, which suggests that patterns could indicate ecosystem weakness under a broad range of environmental conditions. The model independence of our results, provided that they can be reduced to the simplified equation, is also particularly relevant given the diversity of available models and their lack of empirical validation [36]. Our results thus question the prevailing view that self-organized patterns enhance ecosystem resilience [17, 21, 22, 46]. Recent results have also highlighted this possibility, finding that, within relatively limited parameter ranges, self-organized patterns could be transient and unstable, providing a pathway to ecosystem collapse rather than an evasion of it [55].
In summary, our work highlights the importance of isotropy-breaking heterogeneities in determining ecosystem resilience and, more generally, in spatial ecological dynamics. Smaller-scale spatial heterogeneities are thought to lead to more complex spatial patterns [43, 56] and increase ecosystem resilience [54, 13], while we have shown that constant environmental gradients reduce ecosystem resilience, particularly in the presence of vegetation patterns. Future research should combine mathematical modeling with data-driven parameterizations of spatial heterogeneities, considering landscape features, environmental covariates, and anthropogenic pressures [25, 45], to fully characterize how spatial dynamics driven by the intrinsic scales of the water-vegetation feedback interact with endogenous spatial features and determine ecosystem resilience.
Acknowledgments
This work was partially funded by the Center of Advanced Systems Understanding (CASUS), which is financed by Germany’s Federal Ministry of Education and Research (BMBF) and by the Saxon Ministry for Science, Culture and Tourism (SMWK) with tax funds on the basis of the budget approved by the Saxon State Parliament. RMG was also supported by FAPESP through grant ICTP-SAIFR 2021/14335-0.
Supplementary Material
App. A Derivation of the reduced model, Eq. (2.1)
A.1 Interaction-redistribution, kernel-based models.
Within this family of models, we consider an extension of the logistic equation that includes the effect of long-range plant competition and facilitation via nonlocal terms [31, 57],
$\frac{\partial B}{\partial t}=rm_{f}(\tilde{B})\left(1-\frac{B}{K}\right)B-m\,m_{c}(\tilde{B})B+D\frac{\partial^{2}B}{\partial x^{2}},$ (A.1)
where $B$ is the biomass density field, and $x,t$ are the spatial coordinate and time, respectively. $r$ is the intrinsic growth rate in the absence of feedback and stress; $m$ measures the additional mortality due to environmental stress; and $D_{b}$ quantifies vegetation spreading by seed dispersal and/or clonal reproduction. Long-range interactions are introduced by the coefficients $m_{f}(\tilde{B})$ and $m_{c}(\tilde{B})$, which quantify the intensity of facilitation and competition, respectively. Following [57], these coefficients depend both on the weighted average biomass density within a range of a focal location, $\tilde{B}$, and an interaction strength factor $\xi_{f,c}$,
[…]
where $\phi_{f,c}({x^{\prime}})$ are normalized kernels defining the strength of the interaction between vegetation biomass at locations $x$ and $x^{\prime}$. We assume these kernels are normalized Gaussians
$\phi_{f,c}(x)\propto\exp{\left[-\frac{(x-x_{0f,0c})^{2}}{2l_{f,c}^{2}}\right]}$ (A.3)
with standard deviations $l_{f}$ and $l_{c}$ defining the characteristic scale of facilitation and competition, respectively. To account for non-reciprocity in plant interactions, we included a shift $x_{0f,0c}$ so the maximum interaction strength is not achieved at the focal location $x$ and right and left neighbors interact differently with vegetation biomass at $x$. This shift thus breaks the reflection symmetry $x\leftrightarrow-x$ in Eq. (A.1). Note that for vanishing stress and feedback strengths, $m=\xi_{f,c}=0$, Eq. (A.1) reduces to the Fisher-Kolmogorov-Petrovsky-Piskunov (FKPP) equation.
Because Eq. (A.6) is rather complex due to its nonlinearities and nonlocalities, it is hard to work with it analytically and even numerically. We present a method to derive a simpler equation that retains the behavior of the original one near a critical point in parameter space [58, 59]. This method relies on performing a nonlinear change of coordinates to the center manifold of a bifurcation present in the original equation through the following steps. First, we identify a bifurcation in the original equation (our critical point). At that bifurcation, the system has a vanishing eigenvalue, meaning that the dynamics is arbitrarily slow in the direction of the phase space defined by the corresponding eigenvector. This separation of time scales will allow us to reduce the dynamical system to this slow direction because all the other directions in the phase space will rapidly relax and become dynamically irrelevant. Second, we propose a polynomial change of variables using a scaled slow time, assuming that we are close to the critical point. Third, we impose a series of additional restrictions to encapsulate a series of transitions and spatiotemporal scales present in the original models, in particular, a nascent bistability, a Turing instability, and a vanishingly small characteristic wavenumber and group velocity. This set of constraints, together with the original restriction of working close to a bifurcation point, allows us to derive a series of self-consistent reduced equations (independent of the distance to the critical point) that capture increasingly complex phenomena— the more equations we derive, the more complex the phenomena. Lastly, we impose a solvability condition on the hierarchy of relationships obtained in the previous step, which allows us to iteratively obtain the nonlinear change of variables up to a desired order. This iterative procedure progressively refines the accuracy of the reduced equation, although only close to the critical point.
We apply this procedure to the nonlocal model in Eq. (A.1) to obtain the reduced model we used in the main text. First, we write Eq. (A.1) in non-dimensional form by scaling time and the biomass density field
$t=\frac{\tau}{r},\ \ \ \ \ \ \ B=Kb.$ (A.4)
and defining scaled parameters
$\mu=\frac{m}{r}\ \ \ \ \xi_{f,c}=\frac{\chi_{f,c}}{K},\ \ \ \ D=\frac{D_{B}}{r},$ (A.5)
obtaining
$\frac{\partial b}{\partial\tau}=m_{f}(1-b)b-\mu m_{c}b+D\frac{\partial^{2}b}{\partial x^{2}}.$ (A.6)
Next, we find a critical point of Eq. (A.6). To this end, we consider the homogeneous solutions $b_{h}$ satisfying
$0=b_{h}(1-b_{h})e^{\chi_{f}b_{h}}-\mu b_{h}e^{\chi_{c}b_{h}}.$ (A.7)
$b_{h}=0$ is always a solution, and the remaining possible solutions satisfy
$\mu e^{-(\chi_{f}-\chi_{c})b_{h}}=(1-b_{h}).$ (A.8)
Eq. (A.8) defines two curves in the $(b,\mu)$ plane, and the intersections between them correspond to possible equilibria. A bifurcation occurs when the parameters are such that two or more solutions $b_{h}$ collapse to a single point, for example, to $b_{h}=0$. The homogeneous solutions of the nonlocal Eq.  (A.6) present a bifurcation at
$\mu=1$
. Analyzing the linear dynamics,
$b=0+\delta e^{ikx+\lambda t},$
$\lambda=1-\mu-Dk^{2},$ (A.9)
we can identify this bifurcation with a transcritical bifurcation at which the equilibrium $b_{h}=0$ changes its stability from being unstable for $\mu<1$ to stable for $\mu>1$. Having identified a bifurcation, we will state the additional conditions that we will impose:
  1. 1.
    Eq. (A.6) can show bistability or monostability. We will perform our analysis near the transition from one case to the other to cover both.
  2. 2.
    We are also interested in patterns, so we will impose proximity to a Turing instability.
  3. 3.
    Because we are interested in macroscopic patterns, we focus on the case where patterns exhibit big wavelengths, or equivalently, wavenumbers $k\rightarrow 0$.
  4. 4.
    Lastly, the nonreciprocal interaction induces a velocity. We require that the timescale of this movement be the same order of magnitude as that of the evolution of the homogeneous or patterned solution.
These four conditions, together with the proximity to the bifurcation point, reduce our analysis to the neighborhood of a single point in a five-dimensional parameter space (four constraints plus the bifurcation condition). This number of conditions defines the codimension of the reduced equation.
The transition from monostability to bistability can be readily determined by looking for the condition for which, at $b_{h}=0$, the curves defining the remaining equilibria, Eq. (A.8), are tangent to one another. By differentiating (A.8) with respect to $b_{h}$ and evaluating at $b_{h}=0$ for $\mu=1$ one gets
$\chi_{f}-\chi_{c}=1,$
which allows us to obtain the order of magnitude of the biomass as a function of the distance to the bifurcation point. Let
$\mu=1+\epsilon\eta,$
$\chi_{f}-\chi_{c}=1+\chi,$
where $\epsilon\ll 1$ quantifies the distance to the bifurcation, $\eta\sim O(1)$ is a constant, and $\chi\ll 1$ quantifies how far we are from the nascent bistability. Solving for the biomass to the lowest order in Eq. (A.8) leads to
$b_{h}\approx\chi\pm\sqrt{\chi^{2}-2\epsilon\eta}.$
The solution $b_{h}$ is determined from both parameters simultaneously only if $\chi\sim O(\epsilon^{1/2})$. Otherwise, just one of the parameters determines the solution and bistability is lost. This condition leads us to write
$\chi=\epsilon^{1/2}\kappa,$
from which it follows that $b\sim O(\epsilon^{1/2})$.
Next, we proceed to formulate the nonlinear change of variables. We expand the parameters $\mu$, $\chi_{f}$, $\chi_{c}$ and the field $b$ as
$\mu=1+\epsilon\eta+...$
$\chi_{f}-\chi_{c}=1+\epsilon^{1/2}\kappa$
$b=0+B+b(B)^{[2]}+...$
$\dot{B}=a(B)^{[1]}+a(B)^{[2]}+...$
where $(\cdot)^{[n]}$ means terms of polinomial order $n$ in the change of variables coordinate $B$. The linear dynamics (A.9) provides the relevant timescale, $t_{0}=1/\lambda\sim O(\epsilon^{-1})$, and the relevant spatial scale, $l_{0}=1/k\sim O([D/\epsilon]^{1/2})$. Let us for the moment define $(D/\epsilon)^{1/2}=\nu^{-1}$, with $\nu\ll 1$. Then, it follows that the center manifold direction evolves in the slow variables $T=\epsilon\tau$ and $X=\nu x$, and remembering that $b\sim O(\epsilon^{1/2})$, we write $B=\epsilon^{1/2}A(T,X)$. Note that $(db/dt)^{[m]}=a(B)^{[m]}$ by construction. Additionally, $\dot{B}=\epsilon^{3/2}(dA/dT)$, so, to obtain an equation independent of $\epsilon$, upon replacing our change of variables, we consider terms with a prefactor of $\epsilon^{3/2}$ only. Let us compute the first term by replacing our change of variables in Eq. (A.6)
$a(B)^{[1]}=\left(-\epsilon\eta+\nu^{2}D\frac{\partial^{2}}{\partial X^{2}}\right)B.$
For the following terms, we will need to expand the nonlinearities in polynomials, such that
[…]
$=$ $1+\chi_{i}\left(b+c_{1,i}\frac{\partial b}{\partial x}+c_{2,i}\frac{\partial^{2}b}{\partial x^{2}}+...\right)$
$+\frac{\chi_{f}^{2}}{2}\left(b+c_{1,i}\frac{\partial b}{\partial x}+c_{2,i}\frac{\partial^{2}b}{\partial x^{2}}+...\right)^{2}+...$
with
[…]
By collecting the terms quadratic in $B$, we obtain
$a(B)^{[2]}$ $=$ $(\chi_{f}-\chi_{c}-1-\epsilon)B^{2}+\chi_{f}B\left(c_{1f}\frac{\partial B}{\partial x}+c_{2f}\frac{\partial^{2}B}{\partial x^{2}}+...\right)$
$-(1+\epsilon)\chi_{c}B\left(c_{1c}\frac{\partial B}{\partial x}+c_{2c}\frac{\partial^{2}B}{\partial x^{2}}+...\right),$
where not all the derivatives will be relevant because, in the slow variable, they read $(\partial^{n}B/\partial x^{n})=\epsilon^{1/2}\nu^{n}(\partial^{n}A/\partial X^{n})$ with $\nu\sim O(\epsilon^{q})$. We have not specified $q$ yet, but following condition 3 it must be $q>0$. Therefore, at some point in the expansion, the derivatives will have a prefactor with a power of $\epsilon$ greater than $3/2$.
Finally, the cubic term in $B$ reads
$a(B)^{[3]}=-\frac{1}{2}B^{3}+O(\epsilon^{3/2}\nu,\epsilon^{2}).$
We have computed the terms of the nonlinear change of variables $\dot{B}=a(B)^{[1]}+a(B)^{[2]}+a(B)^{[3]}$. Naturally, terms $a(B)^{[4]}$ will be of higher order due to the derived scaling of $b$, $b\sim O(\epsilon^{1/2})$. We explicitly write these terms, replacing the expansion near the critical point $\mu=1+\epsilon\eta$ and $\chi_{f}-\chi_{c}=1+\epsilon^{1/2}\kappa$, obtaining
$\epsilon^{3/2}\frac{\partial A}{\partial T}$ $=$ $\epsilon^{3/2}\left(-\eta A+\kappa A^{2}-\frac{A^{3}}{2}\right)+\epsilon^{1/2}\nu^{2}D\frac{\partial^{2}A}{\partial X^{2}}$
$+\epsilon\nu(\chi_{f}c_{1f}-\chi_{c}c_{1c})A\frac{\partial A}{\partial X}+\epsilon\nu^{2}(\chi_{f}c_{2f}-\chi_{c}c_{2c})A\frac{\partial^{2}A}{\partial X^{2}}$
$+\epsilon\nu^{3}(\chi_{f}c_{3f}-\chi_{c}c_{3c})A\frac{\partial^{3}A}{\partial X^{3}}+\epsilon\nu^{4}(\chi_{f}c_{4f}-\chi_{c}c_{4c})A\frac{\partial^{4}A}{\partial X^{4}}$
$+O(\epsilon^{3/2}\nu,\epsilon^{2},\epsilon\nu^{5}).$
From here, we will impose conditions (2-4) to obtained a closed equation in the limit $\epsilon\rightarrow 0$. The remaining conditions we wish to impose are related to the pattern-formation instability and the induced group velocity. Then, we analyze the linearized equation around a non-zero homogeneous state. Thus, we let $A=A_{h}+\delta e^{ikX+\lambda T}$ and analyze $\lambda(k)$. The imaginary part fulfills
$\epsilon^{3/2}\text{Im}\left[\lambda\right]=\epsilon\nu(\chi_{f}c_{1f}-\chi_{c}c_{1c})kA_{h}-\epsilon\nu^{3}(\chi_{f}c_{3f}-\chi_{c}c_{3c})k^{3}A_{h},$
and the real part
$\epsilon^{3/2}\text{Re}\left[\lambda\right]$ $=$ $\epsilon^{3/2}\left(-\eta+2\kappa A_{h}-\frac{3}{2}A_{h}^{2}\right)-\left(\epsilon^{1/2}\nu^{2}D+\epsilon\nu^{2}A_{h}(\chi_{f}c_{2f}-\chi_{c}c_{2c})\right)k^{2}$
$+\epsilon\nu^{4}A_{h}(\chi_{f}c_{4f}-\chi_{c}c_{4c})k^{4}.$
The Turing instability may occur when the term proportional to $k^{2}$ vanishes. In this limit, the only term stabilizing the equation is the one proportional to $k^{4}$ (provided that it is negative). Then, it follows that
$\epsilon\nu^{4}A_{h}(\chi_{f}c_{4f}-\chi_{c}c_{4c})$ $\sim$ $O(\epsilon^{3/2}),$
$\left(\epsilon^{1/2}\nu^{2}D+\epsilon\nu^{2}A_{h}(\chi_{f}c_{2f}-\chi_{c}c_{2c})\right)$ $=$ $0+O(\epsilon^{3/2}).$ (A.10)
Equations (A.10) are simultaneously solved for $\nu=\epsilon^{1/8}$,
$D=0+\epsilon^{3/4}d,$
and
$c_{2f}+\chi_{c}(c_{2f}-c_{2c})=0+\epsilon^{1/4}\chi_{1}.$
It is easy to verify that the Turing instability occurs for the characteristic wavevector $k_{c}=0+O(\epsilon^{1/8})$, so we are indeed looking at macroscopic patterns of large wavelength consistent with the slow spatial scale $l_{0}$. The remaining condition concerns the induced velocity, which is encapsulated in the imaginary part of $\lambda$. We require that the oscillation of the pattern is on the same timescale as its growth rate. The lowest order corresponds to
$\epsilon^{1+1/8}(\chi_{f}c_{1f}-\chi_{c}c_{1c}),$
Then, it must be fulfilled that
$c_{1f}+\chi_{c}(c_{1f}-c_{1c})=0+s\epsilon^{3/8}.$ (A.11)
Regarding the group velocity induced by the non-reciprocity term, we asuume that, because nonreciprocity is caused by a single process, the nonreciprocal coupling parameters, $x_{0f}$ and $x_{0c}$ have a single origin and are, thus, proportional to a single parameter. Let
$x_{0f}=v\alpha_{f}$
and
$x_{0c}=v\alpha_{c}.$
Then, the Eq. (A.11) is fulfilled for $v=0+\epsilon^{3/8}\alpha$, and it follows that $\epsilon\nu^{3}(\chi_{f}c_{3f}-\chi_{c}c_{3c})\sim\epsilon^{3/2+1/4}$.
Summarizing, considering a region of parameters near the critical point determined by a bifurcation and the four additional conditions we imposed, with the distance to the critical point measured by the bookkeeping parameter $\epsilon\ll 1$, we let
$\mu$ $=$ $1+\epsilon\eta,$
$\chi_{f}$ $=$ $1+\chi_{c}+\epsilon^{1/2}\kappa,$
$\chi_{c}$ $=$ $\frac{l_{f}^{2}}{l_{c}^{2}-l_{f}^{2}}+\epsilon^{1/4}\Gamma,$
$D$ $=$ $0+\epsilon^{3/4}d,$
$v$ $=$ $0+\epsilon^{3/8}\alpha,$
$b$ $=$ $0+\epsilon^{1/2}A(T=\epsilon\tau,X=\epsilon^{1/8}x),$
and insert those expressions in Eq. (A.6). We obtain the equation for $A(X,T)$ reading
$\epsilon^{3/2}\frac{\partial A}{\partial T}$ $=$ $\epsilon^{3/2}\left(-\eta A+\kappa A^{2}-\frac{A^{3}}{2}+(d-(l_{c}^{2}-l_{f}^{2})\Gamma A)\frac{\partial^{2}A}{\partial X^{2}}-3l_{f}^{2}l_{c}^{2}A\frac{\partial^{4}A}{\partial X^{4}}\right.$ (A.13)
$\left.+\alpha\frac{\alpha_{f}l_{c}^{2}-\alpha_{c}l_{f}^{2}}{l_{c}^{2}-l_{f}^{2}}\frac{\partial A}{\partial X}\right)+O(\epsilon^{3/2+1/4}).$
In the limit $\epsilon\rightarrow 0$, this equation describes exactly the behavior of the original system. However, for small values of $\epsilon$, it presents corrections ${}O(\epsilon^{1/4})$ that prevent the application of the reduced equation when we move in the parameter space away from the critical point. Importantly, all the constants in this reduced equation are determined by the ecologically relevant parameters of the starting model. It is easy to verify that one obtains exactly the equation (2.1) in the main text by performing a non-dimensionalization of space. That is, using the nondimensional spatial variable $Z$, defined by $X=(3l_{f}^{2}l_{c}^{2})^{1/4}Z$, and redefining the parameters of spatial interactions accordingly. Explicitly, this non-dimensionalization leads to
$\epsilon^{3/2}\frac{\partial A}{\partial T}$ $=$ $\epsilon^{3/2}\left(-\eta A+\kappa A^{2}-\frac{A^{3}}{2}+\frac{(d-(l_{c}^{2}-l_{f}^{2})\Gamma A)}{(3l_{f}^{2}l_{c}^{2})^{1/2}}\frac{\partial^{2}A}{\partial Z^{2}}-A\frac{\partial^{4}A}{\partial Z^{4}}\right.$
$\left.+\frac{\alpha}{(3l_{f}^{2}l_{c}^{2})^{1/4}}\frac{\alpha_{f}l_{c}^{2}-\alpha_{c}l_{f}^{2}}{l_{c}^{2}-l_{f}^{2}}\frac{\partial A}{\partial Z}\right)+O(\epsilon^{3/2+1/4}).$
A.2 Turing-like coupled water-biomass models
A similar procedure can be performed in models that describe water-vegetation dynamics explicitly using two variables [54, 29, 60]. We consider a classical example of this family of models, originally proposed by Meron and coauthors [61, 62, 47] as a generalization of Klausmeier’s model [32]
$\frac{\partial B}{\partial t}$ $=$ $RBW\left(1-\frac{B}{K}\right)(1+EB)^{2}-MB+D_{B}(B)\frac{\partial^{2}B}{\partial x^{2}},$
$\frac{\partial W}{\partial t}$ $=$ $P-LW-GBW(1+EB)^{2}+D_{W}\frac{\partial^{2}W}{\partial x^{2}}-V\frac{\partial W}{\partial x}.$ (A.14)
where $B$ and $W$ are the biomass and soil water density as a function of space $x$ and time $t$. In the vegetation equation, $R$ is the biomass growth rate per water density, and $M$ and $E$ are the plant mortality rate root-to-shoot ratio, respectively. The diffusion $D_{B}(B)$ accounts for plant dispersal, which we consider non-linear both to account for density-dependent dispersal effects and to ensure that the reduced equation is independent of the expansion parameter $\epsilon$. In the soil water equation, $P$ is the precipitation parameter, $L$ the evaporation rate, and $G$ the water absorption per biomass density rate. $D_{W}$ is the water diffusion rate in the soil, and $V$ modulates the intensity of the water runoff. A range of realistic values for these parameters in a specific ecosystem is provided in [47].
To perform the model reduction, we first assume that $D_{B}=D_{0}+D_{1}B+...$ and define the following dimensionless variables and parameters,
$t$ $=\frac{\tau}{L},$ $x$ $=\sqrt{\frac{D_{W}}{L}}z,$ $B$ $=Kb,$
$W$ $=\frac{L}{R}w,$ $p$ $=\frac{PR}{L^{2}},$ $m$ $=\frac{M}{L},$
$\delta$ $=EK,$ $d_{0}$ $=\frac{D_{0}}{D_{W}},$ $d_{1}$ $=\frac{D_{1}K}{D_{W}},$
$\gamma$ $=GK,$ $s$ $=\frac{V}{\sqrt{LD_{W}}}.$
Using these new quantities, we can rewrite Eqs. (A.14) as
$\frac{\partial}{\partial\tau}\begin{pmatrix}b\\ w\end{pmatrix}=\begin{pmatrix}bw(1-b)(1+\delta b)^{2}-mb\\ p-w-\gamma bw(1+\delta b)^{2}\end{pmatrix}+\begin{pmatrix}(d_{0}+d_{1}b)\frac{\partial^{2}b}{\partial z^{2}}\\ \frac{\partial^{2}u}{\partial z^{2}}-s\frac{\partial u}{\partial z}\end{pmatrix}.$
whose homogeneous solutions satisfy
$w_{h}{b_{h}(1-b_{h})(1+\delta b_{h})^{2}}$ $=$ ${mb_{h}},$
$w_{h}$ $=$ $\frac{p}{1+\gamma b(1+\delta b)^{2}}.$
One solution corresponds to the trivial bare-soil solution
$\begin{pmatrix}b_{0}\\ w_{0}\end{pmatrix}=\begin{pmatrix}0\\ p\end{pmatrix}.$
and the nontrivial solutions correspond to the intersection of the curves
$w_{b}(b)$ $=$ $\frac{m}{(1-b_{h})(1+\delta b_{h})^{2}},$
$w_{w}(b)$ $=$ $\frac{p}{1+\gamma b(1+\delta b_{h})^{2}},$
in the plane $(w,b)$. We find the nascent bistability by imposing that these equilibria collapse to a single point simultaneously. This is achieved by imposing that
$\left.\frac{dw_{b}}{db_{h}}\right|_{b_{h}=0}$ $=$ $\left.\frac{w_{w}}{db_{h}}\right|_{b_{h}=0},$
$-m\left(2\delta-1\right)$ $=$ $-p\gamma.$
As we did for the nonlocal model, we further impose this condition at a bifurcation of the $b_{0}=0$ state. We additionally impose the existence of a Turing instability, that spatial structures are macroscopic ($k\rightarrow 0$), and that the velocity of these structures is on the same timescale as the growth-rate of homogeneous and patterned solutions (basically, the same conditions 1-4 mentioned in the previous section).
To derive the reduced equation, we write the model using the bare soil solution as the reference state,
$w=p+u,$
and expand the nonlinearities in polynomials. After these manipulations, the system reads
$\frac{\partial}{\partial\tau}\begin{pmatrix}b\\ u\end{pmatrix}=\begin{pmatrix}p-m&0\\ -\gamma p&-1\end{pmatrix}\begin{pmatrix}b\\ u\end{pmatrix}+\begin{pmatrix}f(b,u)+(d_{0}+d_{1}b)\frac{\partial^{2}b}{\partial z^{2}}\\ -\gamma g(b,u)+\frac{\partial^{2}u}{\partial z^{2}}-s\frac{\partial u}{\partial z}\end{pmatrix},$ (A.15)
where
$f(b,u)$ $=$ $ub+b^{2}(2\delta p-p)+b^{3}(p\delta^{2}-2p\delta)+ub^{2}(2\delta-1)+O(b^{4},ub^{3}),$ (A.16)
$g(b,u)$ $=$ $ub+b^{2}2\delta p+b^{3}p\delta^{2}+ub^{2}2\delta+O(ub^{3}).$ (A.17)
The linear part of Eq. (A.15) shows a change in the stability of the bare soil $(b,\,u)=(0,0)$ when $p=m$, and the form of the linear part (the Jacobian), one deduces it is a transcritical bifurcation. Letting
$p=m-\epsilon\eta,$
the slow eigenvalue is
$\lambda=-\epsilon\eta.$
At the bifurcation point $\epsilon=0$, the tangent to the center manifold is the eigenvector of the Jacobian
$\mathbf{J}=\begin{pmatrix}0&0\\ -\gamma p&-1\end{pmatrix},$
which is
$\mathbf{v}=C\begin{pmatrix}1\\ -\gamma p\end{pmatrix}.$
This is our starting point to obtain the change of variables to the center manifold in the neighborhood of the bifurcation and our further conditions. The amplitude along this direction in phase space is the variable we will describe; let us call it $C$. This amplitude will evolve spatio-temporally on a slow timescale—because we are close to a bifurcation—and on a slow spatial scale—because we will impose that the wavevector $k\rightarrow 0$. Let these scales correspond to $t_{0}=1/\lambda\sim O(\epsilon^{-1})$ and $l_{0}=1/k\sim O(\nu^{-1})$, with $\nu$ related to $\epsilon$ but unknown for the moment. We define the spatio-temporal slow variables $T=\epsilon\tau$ and $Z=\nu z$. We note that the nascent bistability condition evaluated at the bifurcation point reads
$\gamma=2\delta-1.$
Again, solving for the homogeneous solutions at lowest order reveals that for bistability to exist, it must be satisfied that
$\gamma=2\delta-1-\epsilon^{1/2}\kappa,$
from which it follows that $b_{h}\sim O(\epsilon^{1/2})$. With those ingredients, we can write for Eq. (A.15) the following
$\epsilon\frac{\partial}{\partial T}\begin{pmatrix}b\\ u\end{pmatrix}=\mathbf{J}\begin{pmatrix}b\\ u\end{pmatrix}+\begin{pmatrix}f(b,u)\\ -\gamma g(b,u)\end{pmatrix}+\begin{pmatrix}-\epsilon\eta b\\ 0\end{pmatrix}+\begin{pmatrix}\nu^{2}(d_{0}+d_{1}b)\frac{\partial^{2}b}{\partial Z^{2}}\\ \nu^{2}\frac{\partial^{2}u}{\partial Z^{2}}-\nu s\frac{\partial u}{\partial Z}\end{pmatrix},$ (A.18)
and perform the change of variables such that
$\begin{pmatrix}b\\ u\end{pmatrix}$ $=$ $\epsilon^{1/2}A\mathbf{v}+\epsilon^{1/2}\nu\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,1]}+\epsilon\nu\begin{pmatrix}b(C)\\ u(C)\end{pmatrix}^{[1,1]}+...,$
$\epsilon^{3/2}\frac{\partial A}{\partial T}$ $=$ $\epsilon^{1/2}a(A)^{[1/2,0]}+\epsilon^{1/2}\nu a(A)^{[1/2,1]}+\epsilon\nu a(A)^{[1,1]}+....$ (A.19)
$(\cdot)^{[n,m]}$ are just bookkeeping superscripts to label the functions proportional to $\epsilon^{n}$ and $\nu^{m}$. We insert those expressions into Eq. (A.18) and solve iteratively the resulting hierarchy, noting that
$\epsilon\frac{\partial}{\partial\tau}\begin{pmatrix}b\\ u\end{pmatrix}^{[n,m]}=\epsilon^{n}\nu^{m}a(A)^{[n,m]}\mathbf{v}.$
The first-order equation in this hierarchy corresponds to the linearized dynamics, reading
$O(\epsilon^{1/2}\nu^{0}):$
$\epsilon^{1/2}a(A)^{[1/2,0]}\mathbf{v}=\epsilon^{1/2}A\mathbf{J}\mathbf{v},$
and considering how $\mathbf{J}$ acts on $\mathbf{v}$, we find that $a(A)^{[1/2,0]}$ vanishes. We solve iteratively for the next order
$O(\epsilon^{1/2}\nu^{1}):$
$\epsilon^{1/2}\nu a(A)^{[1/2,1]}\mathbf{v}=\epsilon^{1/2}\nu\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,1]}+\epsilon^{1/2}\nu\begin{pmatrix}0\\ \gamma ps\frac{\partial A}{\partial Z},\end{pmatrix}$
which has to be solved for two unknowns, $a(A)^{[1/2,1]}$ and the vector of the change of variables $\left((b(A),u(A))^{[1/2,1]}\right)^{T}$ (where $T$ means the transpose).
Note that the problem is linear, so solutions must fulfill the solvability condition. The solvability condition states that the system $\mathbf{A}\mathbf{x}=\mathbf{b}$ has a solution if and only if $\mathbf{b}\perp\text{Ker}\{\mathbf{A}^{\dagger}\}$, where $(\cdot)^{\dagger}$ means the adjoint, or conjugated transpose under the usual inner product. Therefore, we need to compute the kernel of the operator
$\mathbf{J}^{\dagger}=\begin{pmatrix}0&-\gamma p\\ 0&-1\end{pmatrix}.$
which corresponds to
$\text{Ker}\{\mathbf{J}^{\dagger}\}=\left\{\begin{pmatrix}1\\ 0\end{pmatrix}\right\}.$
Next, we impose that
$\left[a(A)^{[1/2,1]}\mathbf{v}-\begin{pmatrix}0\\ \gamma ps\frac{\partial A}{\partial Z}\end{pmatrix}\right]\perp\begin{pmatrix}1\\ 0\end{pmatrix},$
or equivalently
$\left[a(A)^{[1/2,1]}\mathbf{v}-\begin{pmatrix}0\\ \gamma ps\frac{\partial A}{\partial Z}\end{pmatrix}\right]\cdot\begin{pmatrix}1\\ 0\end{pmatrix}=0.$
This implies that
$a(A)^{[1/2,1]}=0,$
which leads to the equation
$\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,1]}=-\begin{pmatrix}0\\ \gamma ps\frac{\partial A}{\partial Z}\end{pmatrix}.$
This equation has multiple solutions. Nevertheless, the method warrants that the normal form (or the reduced equation) is unique. Then, one must choose the solution such that the change of variables (A.19) cannot be simplified further with additional changes of variables. We choose
$\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,1]}=\begin{pmatrix}0\\ \gamma ps\frac{\partial A}{\partial Z}\end{pmatrix}.$
The next order reads
$O(\epsilon^{1}\nu^{0}):$
$\epsilon a(A)^{[1,0]}\mathbf{v}=\epsilon\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1,0]}+\epsilon\begin{pmatrix}0\\ -\gamma pA^{2}\end{pmatrix}.$
Note that we face the same linear problem as before, this is what makes the method iterable up to the desired order. Then, applying the same arguments as in the previos order, we compute the solutions
$a(A)^{[1,0]}$ $=$ $0,$
$\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1,0]}$ $=$ $\begin{pmatrix}0\\ -\gamma pA^{2}\end{pmatrix}.$
That leads to the next order
$O(\epsilon^{1/2}\nu^{2}):$
$\epsilon^{1/2}\nu^{2}a(A)^{[1/2,2]}\mathbf{v}=\epsilon^{1/2}\nu^{2}\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,2]}+\epsilon^{1/2}\nu^{2}\begin{pmatrix}d_{0}\\ -\gamma p-\gamma ps^{2}\end{pmatrix}\frac{\partial^{2}A}{\partial Z^{2}},$
with solutions
$a(A)^{[1/2,2]}$ $=$ $d_{0}\frac{\partial^{2}A}{\partial Z^{2}},$
$\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,2]}$ $=$ $\begin{pmatrix}0\\ -(d_{0}+\gamma p+\gamma ps^{2})\frac{\partial^{2}A}{\partial Z^{2}}\end{pmatrix}.$
At the next order, we find
$O(\epsilon^{1/2}\nu^{3}):$
$\epsilon^{1/2}\nu^{3}a(A)^{[1/2,3]}\mathbf{v}=\epsilon^{1/2}\nu^{3}\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,3]}+\epsilon^{1/2}\nu^{3}\begin{pmatrix}0\\ \gamma ps+s(d_{0}+\gamma p+\gamma ps^{2})\end{pmatrix}\frac{\partial^{3}A}{\partial Z^{3}},$
with solutions
$a(A)^{[1/2,3]}$ $=$ $0,$
$\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,3]}$ $=$ $\begin{pmatrix}0\\ [\gamma ps+s(d_{0}+\gamma p+\gamma ps^{2})]\frac{\partial^{3}A}{\partial Z^{3}}\end{pmatrix}.$
Next
$O(\epsilon^{1/2}\nu^{4}):$
$\epsilon^{1/2}\nu^{4}a(A)^{[1/2,4]}\mathbf{v}=\epsilon^{1/2}\nu^{4}\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,4]}+\epsilon^{1/2}\nu^{4}\begin{pmatrix}0\\ -\gamma ps^{2}-s^{2}(d_{0}+\gamma p+\gamma ps^{2})-(d_{0}+\gamma p+\gamma ps^{2})\end{pmatrix}\frac{\partial^{4}A}{\partial Z^{4}},$
with solutions
$a(A)^{[1/2,4]}$ $=$ $0,$
$\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1/2,4]}$ $=$ $\begin{pmatrix}0\\ -\gamma ps^{2}-s^{2}(d_{0}+\gamma p+\gamma ps^{2})-(d_{0}+\gamma p+\gamma ps^{2})\end{pmatrix}\frac{\partial^{4}A}{\partial Z^{4}}.$
Next
$O(\epsilon^{1}\nu^{1}):$
$\epsilon\nu a(A)^{[1,1]}\mathbf{v}=\epsilon\nu\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1,1]}+\epsilon\nu\begin{pmatrix}\gamma psA\frac{\partial A}{\partial Z}\\ -\gamma^{2}psA\frac{\partial A}{\partial Z}+2\gamma psA\frac{\partial A}{\partial Z}\end{pmatrix},$
with solutions
$a(A)^{[1,1]}$ $=$ $\gamma psA\frac{\partial A}{\partial Z},$
$\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1,1]}$ $=$ $\begin{pmatrix}0\\ \gamma^{2}p^{2}s-\gamma^{2}ps+2\gamma ps\end{pmatrix}A\frac{\partial A}{\partial Z}.$
Next
$O(\epsilon^{1}\nu^{2}):$
$\epsilon\nu^{2}a(A)^{[1,2]}\mathbf{v}=\epsilon\nu^{2}\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1,2]}+\epsilon\nu^{2}\begin{pmatrix}[-(d_{0}+\gamma p+\gamma ps^{2})+d_{1}]A\frac{\partial^{2}A}{\partial Z^{2}}\\ \gamma(d_{0}+\gamma p+\gamma ps)A\frac{\partial^{2}A}{\partial Z^{2}}-\gamma p\frac{\partial^{2}A^{2}}{\partial Z^{2}}-s(\gamma^{2}p^{2}s-\gamma^{2}ps+2\gamma ps)\frac{\partial(A\frac{\partial A}{\partial Z})}{\partial Z}\end{pmatrix},$
with solutions
$a(A)^{[1,2]}$ $=$ $[-(d_{0}+\gamma p+\gamma ps^{2})+d_{1}]A\frac{\partial^{2}A}{\partial Z^{2}},$
$\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1,2]}$ $=$ $\begin{pmatrix}0\\ \left(-\gamma p[-(d_{0}+\gamma p+\gamma ps^{2})+d_{1}]A\frac{\partial^{2}A}{\partial Z^{2}}+\gamma(d_{0}+\gamma p+\gamma ps)A\frac{\partial^{2}A}{\partial Z^{2}}-\right.\\ \left.\gamma p\frac{\partial^{2}A^{2}}{\partial Z^{2}}-s(\gamma^{2}p^{2}s-\gamma^{2}ps+2\gamma ps)\frac{\partial(A\frac{\partial A}{\partial Z})}{\partial Z}\right)\end{pmatrix}.$
Next
$O(\epsilon^{1}\nu^{3}):$
$\epsilon\nu^{3}a(A)^{[1,3]}\mathbf{v}=\epsilon\nu^{3}\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1,3]}+\epsilon\nu^{3}\begin{pmatrix}[\gamma ps+s(d_{0}+\gamma p+\gamma ps^{2})]A\frac{\partial^{3}A}{\partial Z^{3}}\\ -\gamma[\gamma ps+s(d_{0}+\gamma p+\gamma ps^{2})]A\frac{\partial^{3}A}{\partial Z^{3}}+\frac{\partial^{2}u(A)^{[1,1]}}{\partial Z^{2}}-s\frac{\partial u(A)^{[1,2]}}{\partial Z}\end{pmatrix},$
where the terms $u(A)^{[1,1]}$ and $u(A)^{[1,2]}$ have been previously computed and are cumbersome to write (nevertheless, they will have no impact at the end). The solution is
$a(A)^{[1,3]}$ $=$ $[\gamma ps+s(d_{0}+\gamma p+\gamma ps^{2})]A\frac{\partial^{3}A}{\partial Z^{3}},$
$\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1,3]}$ $=$ $\begin{pmatrix}0\\ \left(-\gamma p[\gamma ps+s(d_{0}+\gamma p+\gamma ps^{2})]A\frac{\partial^{3}A}{\partial Z^{3}}-\right.\\ \left.\gamma[\gamma ps+s(d_{0}+\gamma p+\gamma ps^{2})]A\frac{\partial^{3}A}{\partial Z^{3}}+\frac{\partial^{2}u(A)^{[1,1]}}{\partial Z^{2}}-s\frac{\partial u(A)^{[1,2]}}{\partial Z}\right)\end{pmatrix}.$
Next
$O(\epsilon^{1}\nu^{4}):$
$\epsilon\nu^{4}a(A)^{[1,4]}\mathbf{v}=\epsilon\nu^{4}\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[1,4]}+\epsilon\nu^{4}\begin{pmatrix}[-\gamma ps^{2}-s^{2}(d_{0}+\gamma p+\gamma ps^{2})-(d_{0}+\gamma p+\gamma ps^{2})]A\frac{\partial^{4}A}{\partial Z^{4}}\\ ...\end{pmatrix},$
We can get the solution for the dynamics of the center manifold coordinate
$a(A)^{[1,4]}=[-\gamma ps^{2}-s^{2}(d_{0}+\gamma p+\gamma ps^{2})-(d_{0}+\gamma p+\gamma ps^{2})]A\frac{\partial^{4}A}{\partial Z^{4}},$
The correction to the change of variables will not be needed, thus we do not write it explicitly. The last relevant contribution (we will check this a posteriori) will be the next order
$O(\epsilon^{3/2}\nu^{0}):$
$\epsilon^{3/2}a(A)^{[3/2,0]}\mathbf{v}=\epsilon^{3/2}\mathbf{J}\begin{pmatrix}b(A)\\ u(A)\end{pmatrix}^{[3/2,0]}+\epsilon^{3/2}\begin{pmatrix}-\eta A+\kappa pA^{2}+p(4\delta-3\delta^{2})A^{3}\\ ...\end{pmatrix},$
and the correction to the dynamics
$a(A)^{[3/2,0]}=-\eta A+\kappa pA^{2}+p(4\delta-3\delta^{2})A^{3}.$
So far, we have found that
$\epsilon^{3/2}\frac{\partial A}{\partial T}$ $=$ $\epsilon^{3/2}\left(-\eta A+\kappa pA^{2}+p(4\delta-3\delta^{2})A^{3}\right)+\epsilon^{1/2}\nu^{2}d_{0}\frac{\partial^{2}A}{\partial Z^{2}}+\epsilon\nu\gamma psA\frac{\partial A}{\partial Z}$
$+\epsilon\nu^{2}[-(d_{0}+\gamma p+\gamma ps^{2})+d_{1}]A\frac{\partial^{2}A}{\partial Z^{2}}+\epsilon\nu^{3}[\gamma ps+s(d_{0}+\gamma p+\gamma ps^{2})]A\frac{\partial^{3}A}{\partial Z^{3}}$
$+\epsilon\nu^{4}[-\gamma ps^{2}-s^{2}(d_{0}+\gamma p+\gamma ps^{2})-(d_{0}+\gamma p+\gamma ps^{2})]A\frac{\partial^{4}A}{\partial Z^{4}}+O(\epsilon^{3/2}\nu,\epsilon\nu^{5},\epsilon^{2}).$
Then, similarly as the case of the nonlocal model, one imposes the conditions for the Turing instability at vanishing wavenumber and the oscilation frequency to be of the same order as the grow rate. One finds that the conditions are satisfied for
$\nu$ $=$ $\epsilon^{1/8},$
$d_{0}$ $=$ $\epsilon^{3/4}d,$
$s$ $=$ $\epsilon^{3/8}\alpha,$
$d_{1}$ $=$ $\gamma p-\epsilon^{1/4}\Gamma.$
Finally, one finds the equation
$\epsilon^{3/2}\frac{\partial A}{\partial T}$ $=$ $\epsilon^{3/2}\left(-\eta A+p\kappa A^{2}+p(4\delta-3\delta^{2})A^{3}+d\frac{\partial^{2}A}{\partial Z^{2}}\right.$ (A.20)
$\left.+\alpha\gamma pA\frac{\partial A}{\partial Z}-\Gamma A\frac{\partial^{2}A}{\partial Z^{2}}-\gamma pA\frac{\partial^{4}A}{\partial Z^{4}}\right)+O(\epsilon^{3/2+1/8}).$
Summarizing, the critical conditions and the change of variables are as follows
$p$ $=$ $m-\epsilon\eta,$
$\gamma$ $=$ $2\delta-1-\epsilon^{1/2}{\kappa},$
$d_{1}$ $=$ $\gamma p-\epsilon^{1/4}\Gamma,$
$s$ $=$ $\epsilon^{3/8}\alpha,$
$d_{0}$ $=$ $\epsilon^{3/4}d,$
$\begin{pmatrix}b\\ u\end{pmatrix}$ $=$ $\epsilon^{1/2}A(T=\epsilon t,Z=\epsilon^{1/8}z)\begin{pmatrix}1\\ -\gamma p\end{pmatrix}+\begin{pmatrix}0\\ -\epsilon^{3/4}\gamma p\frac{\partial^{2}A}{\partial Z^{2}}-\epsilon\gamma pA^{2}-\epsilon\gamma p\frac{\partial^{4}A}{\partial Z^{4}}+\epsilon\gamma p\alpha\frac{\partial A}{\partial Z}\end{pmatrix},$
inserting them in Eq. (A.18), one finds Eq. (A.20), which correspods to the main text equation up to a reescaling of the spatial coordinate and a redefinition of the corresponding parameters. Explicitly, one lets $A=H/[2p|4\delta-3\delta^{2}|]^{1/2}$ and $Z=Y(\gamma p)^{1/4}/[2p|4\delta-3\delta^{2}|]^{1/8}$, obtaining
$\epsilon^{3/2}\frac{\partial H}{\partial T}$ $=$ $\epsilon^{3/2}\left(-\eta H+\frac{p\kappa}{[2p|4\delta-3\delta^{2}|]^{1/2}}H^{2}-\frac{1}{2}H^{3}+\frac{d[2p|4\delta-3\delta^{2}|]^{1/4}}{(\gamma p)^{1/2}}\frac{\partial^{2}H}{\partial Y^{2}}\right.$ (A.21)
$\left.+\frac{\alpha\gamma p[2p|4\delta-3\delta^{2}|]^{1/8}}{(\gamma p)^{1/4}}H\frac{\partial H}{\partial Y}-\frac{\Gamma[2p|4\delta-3\delta^{2}|]^{1/4}}{(\gamma p)^{1/2}}H\frac{\partial^{2}H}{\partial H^{2}}-H\frac{\partial^{4}H}{\partial Y^{4}}\right)+O(\epsilon^{3/2+1/8}),$
which corresponds exactly to the main text equation. The corrections are of the $\epsilon^{1/8}$ order when away from the bifurcation point, setting the region of validity of the reduced equation.
App. B Numerical method
We performed all the numerical simulations using finite differences for the spatial discretization ($dx=0.5$) and the fifth-order Dormand–Prince method with adaptative time step (DOPRI5) for the time integration. For each set of parameters ($\Gamma,\alpha$), we computed the bifurcation diagrams initializing the simulations with favorable conditions (low environmental stress, $\eta\ll 0$), ensuring that the system is in the quasi-homogeneous solution. Then, we recursively increased $\eta$ in small steps ($\sim 0.0026$) quasistatically, that is, allowing the system to reach the new equilibrium state between two consecutive variations in $\eta$. We defined the equilibrium state using a tolerance criterion, such that for every point of the simulation, it held that $b_{t+dt}-b_{t}<dt\cdot T$, with $T=10^{-5}$. We ran a sensitivity analysis on the value of $T$ and did not find significant changes in our results using lower values of $T$.
App. C The phenomenon in the original models
For completeness, we performed numerical simulations of the original models we used to derive the reduced equation (Fig. 5). Panels a) and b) show typical bifurcation diagrams for the stability of the vegetated solution in the reaction-diffusion and non-local interaction-redistribution model. In both panels, square markers correspond to non-pattern forming cases (homogeneous vegetation covers) and circles correspond to pattern-forming cases. Increasing the parameter responsible for nonreciprocal interactions—$x_{0c}$ in the nonlocal model, and $s$ in the water-biomass one—the vegetation in the bulk dies at levels of environmental stress—$\mu$ in the non-local model and $m$ in the water-biomass—the ecosystems with reciprocal interactions can withstand. Panels c) and d) show the value of the environmental stress at which a convective instability arises relative to the tipping point in the non-spatial limit. We compute this environmental threshold for different values of the parameter responsible for the pattern formation instability—$l_{c}$ in the nonlocal model and $d_{0}$ in the water-biomass—and the non-reciprocity parameter—$x_{0c}$ in the nonlocal model and $s$ in the water-biomass models. These numerical simulations in the original models confirm the two key results we obtained with the reduced equation. First, nonreciprocal interactions reduced ecosystem resilience for a broad range of parameters. Second, ecosystems exhibiting spatial patterns are more sensitive to convective instabilities than their homogeneous counterparts. That is, patterned ecosystems collapse due to convective instabilities at levels of environmental stress that homogeneous vegetation covers can withstand.
Figure 5: Effect of convective instabilities on the ecosystem resilience using the nonlocal (A, C) and water-vegetation (B, D) models we used to derive the reduced equation. A) Bifurcation diagram for the nonlocal model. B) Bifurcation diagram in the water-biomass model. In both panels, square markers indicate parameters for which a Turing instability does not occur, circle markers indicate parameters for which a patterned solution is observed. The color in each curve indicates the distance between the tipping point of the non-spatial model and the maximum environmental stress that ecosystems with the curve-specific parameterization can withstand. Blue (red) indicates an increased (reduced) resilience relative to the non-spatial case. C, D) Environmental stress level at which convective instabilities arise (relative to the non-spatial tipping point) as a function of two parameters: one controlling the pattern formation instability—$l_{c}$ and $d_{0}$ for the nonlocal and water-biomass model, respectively—and the other measuring the intensity of nonreciprocal interactions —-$x_{0c}$ and $s$ in the nonlocal and water-biomass model, respectively. The nonlocal model parameters were $\chi_{f}=3.3$, $\chi_{c}=2$, $D=0.3$, $l_{f}=1$, and $x_{0f}=0$. The water-biomass model parameters were $p=1$, $\delta=1$, $\gamma=0.5$, and $d_{1}=0.001$.